Wide Field Camera 3
This webpage is kept for archival purposes only and is no longer updated or maintained.

Project Updates 1999


Eighth SOC meeting held at KSC. New budgetary threats discussed.

Three instrument upgrade CCRs approved by Project and SOC.

IR detector flight multiplexor planning complete. PDR scheduled for January.

CCD detector proposals received from vendors.

HST SM3A completed.


Instrument descope plan started.

Characterization of IR detectors complete. The detectors meet minimum requirements, discussing possible improvements for flight. Flight detector procurement under way.

IR enclosure thermal testing successfully completed at Ball. The enclosure design meets expectations for 150K operation.

The above two tasks complete the technical feasibility demonstrations for the IR channel.


The SOC held its seventh meeting at STScI.

Status of the 10 major "upgrades" to the instrument were reviewed by the SOC. Work has narrowed down to 5 of the 10. Resolution planned by December 1, in time for the next SOC meeting at KSC.

Initial IR detector characterization completed on Rockwell devices 101 and 103. These trailblazer devices with a 1.7 um cutoff met expectations.

Budget issues are being resolved as the NASA budget becomes more certain.


SRR RFAs completed.

Optical bench thermal concept complete (includes cold bench operation).

EEV CCD test devices arrive at STScI and delivered to GSFC.

Significant progress in the SITe/UofA effort for the CCD coatings.

Preliminary HgCdTe tests completed by the DCL (two devices).

Recent budget activities encouraged some budget exercises and preliminary descope planning.


A very successful SRR was completed.

SRR Requests for Action (RFAs) being worked.

A formal plan for disposition of the Configuration Change Request (CCR) upgrades to the instrument was agreed to by management.

CCRs being worked.

HgCdTe detector characterization started in the DCL.

Stimulus requirements and current capabilities being analyzed.

EEV CCD procurement plans changed by management.


A very successful Filter Workshop was held at STScI.

The SOC met immediately afterwards and constructed the filter list for the UVIS and IR channels.

Definition begun for optical bench temperature (driven by the IR channel background).

EEV CCD formal procurement activities started.


Testing of Site/UofA 1100x330 CCD devices begun.

Rockwell 1K x 1K HgCdTe demonstration detectors completed.

Preparations for the August System Requirements Review begun.

Completed detailed documentation of the trade studies.


Preliminary designs for the integrated, dual-channel WFC3 completed. Detailed design and packaging is continuing.

Filter workshop planning completed. The workshop will be in July.

The WFC3 Contract End-Item Specification is in final review.

The WFC3 Web-based document management system is on-line. A new server for WFC3 Web and data is online.


WF/PC (1) metrology and deintegration completed.

Trade studies completed for the IR channel - 1K x 1K detector format selected.

Detailed design studies continuing.

SOC meeting at STScI.


The Origins Subcommittee Report from its late February meeting strongly endorses the scientific case for a near-IR channel for WFC3 based on a preliminary concept from the Project. Detailed studies begin on the two channel camera.

Discussions begun with EEV for potential WFC3 CCD detectors.


Scientific Oversight Committee (SOC) meeting at GSFC.

CCD procurement delayed because of administrative changes at Lockheed Martin Fairchild Systems in Tustin.

Inventory for WF/PC (1) residual hardware completed.


All optical elements deintegrated from the WF/PC (1) instrument. The optical bench, latches, and radiator are the major remaining elements to deintegrate.

System-level trades and design begins, assuming CCD and IR channels.

Lot 8 Site devices delayed.